### v1

# OFFSIDE: Benchmarking Unlearning Misinformation in Multimodal Large Language Models

Abstract
Advances in Multimodal Large Language Models (MLLMs) intensify concerns about data privacy, making Machine Unlearning (MU), the selective removal of learned information, a critical necessity. However, existing MU benchmarks for MLLMs are limited by a lack of image diversity, potential inaccuracies, and insufficient evaluation scenarios, which fail to capture the complexity of real-world applications. To facilitate the development of MLLMs unlearning and alleviate the aforementioned limitations, we introduce OFFSIDE, a novel benchmark for evaluating misinformation unlearning in MLLMs based on football transfer rumors. This manually curated dataset contains 15.68K records for 80 players, providing a comprehensive framework with four test sets to assess forgetting efficacy, generalization, utility, and robustness. OFFSIDE supports advanced settings like selective unlearning and corrective relearning, and crucially, unimodal unlearning (forgetting only text data). Our extensive evaluation of multiple baselines reveals key findings: (1) Unimodal methods (erasing text-based knowledge) fail on multimodal rumors; (2) Unlearning efficacy is largely driven by catastrophic forgetting; (3) All methods struggle with "visual rumors" (rumors appear in the image); (4) The unlearned rumors can be easily recovered and (5) All methods are vulnerable to prompt attacks. These results expose significant vulnerabilities in current approaches, highlighting the need for more robust multimodal unlearning solutions. The code is available at https://github.com/zh121800/OFFSIDE.
1 Introduction
With the rapid development and widespread application of multimodal large language models (MLLMs), models pre-trained on large-scale corpora can quickly adapt to various downstream tasks, such as visual question answering (Antol et al., 2015; Goyal et al., 2017; Pang et al., 2025), visual understanding (Sugiyama et al., 2007; Guo et al., 2016; Li et al., 2024c; Xu et al., 2025a), and reasoning (Johnson et al., 2017; Perez et al., 2018; Li et al., 2025a; Wang et al., 2025). However, during both the pretraining and post-training phases, unwanted content, such as private information and rumors, may be included, which could lead to the leakage of personal privacy and the spread of misinformation. These raise concerns about the security of MLLMs (Chen et al., 2025). Machine Unlearning (MU) (Wang et al., 2024b; Deng et al., 2025) has been proposed to address these ethical and security concerns in MLLMs, aiming to eliminate the influence of unwanted data and its effects on model performance without requiring retraining from scratch, while also complying with legal frameworks (Dang, 2021).
Given that MLLMs integrate knowledge across multiple modalities, a growing line of work has begun to study MU within multimodal contexts (Liu et al., 2024c; Xu et al., 2025b; Dontsov et al., 2024; Li et al., 2024b). However, existing benchmarks commonly rely on generative models (e.g., Arc2Face (Papantoniou et al., 2024), Flux (Labs, 2024), StyleGAN2 (Karras et al., 2020), and DuelGAN (Wei et al., 2022)) to synthesize images, neglecting harmful cues embedded in the visual modality and risking the introduction of biases that diverge from real-world distributions (Westerlund, 2019; Dolhansky et al., 2020). Moreover, existing benchmarks fail to support selectively removing specific information in an image while preserving unrelated information, typically deleting all text linked to a given image (Cheng et al., 2023). In addition, they pay little attention to the downstream effects of unlearning on other post-training procedures, such as continual learning (Wang et al., 2024a), despite its importance for evaluating practical utility  (Van de Ven & Tolias, 2019). Taken together, these limitations result in an incomplete assessment of multimodal unlearning, underscoring the need for a comprehensive benchmark tailored to MLLMs.
Table 1: Benchmark Comparison. OFFSIDE is the first to support (1) multi-image entity association (group images for each player), (2) selective unlearning of private attributes while preserving shared knowledge, (3) corrective relearning, a continual learning setting, and (4) Unimodal unlearning (unlearn through only pure text data).
Benchmark Text Image Setting
Type Source Entity Complete Selective Corrective Unimodal
Association Unlearning Unlearning relearning Unlearning
MUSE (Shi et al., 2024) - - -
TOFU (Maini et al., 2024) - - -
MMUBench (Li et al., 2024b) Real World MIKE (Li et al., 2024a) multiple
MLLMU-Bench (Liu et al., 2024c) Synthetic Arc2Face (Papantoniou et al., 2024) Single
PEBench (Xu et al., 2025b) Synthetic Flux (Labs, 2024) multiple
CLEAR (Dontsov et al., 2024) Synthetic StyleGAN2 (Karras et al., 2020) multiple
OFFSIDE (Ours) Real World Google multiple
In this view, we propose OFFSIDE, a benchmark inspired by football transfer market rumors, aimed at simulating diverse real-world scenarios. OFFSIDE consists of 15.68K manually created Vision-Question-Answer (VQA) pairs, with 7.84K dedicated to the multimodal unlearning and 7.84K for the unimodal unlearning. It features four distinct datasets: Forget Set, Retain Set, Test set and Relearn Set, each designed to evaluate specific aspects of unlearning methods, including unlearning efficacy, generalizability, model utility, and robustness, across both multimodal and unimodal settings. A comprehensive comparison between previous benchmarks and OFFSIDE is shown in Table 1. In OFFSIDE, each player (entity) is linked to a set of images containing both private information (e.g., transfer records) and shared information (e.g., age, height, and name). The diverse text-image connections are designed for the selective unlearning setting, which only removes the private information of the target rumor and saves the shared ones. Meanwhile, we also simulate a corrective relearning framework to investigate the sources of the unlearning abilities of the tested methods.
As shown in Figure 1. We define four real-world scenarios for evaluation: the Complete Unlearning: setting tests whether unlearning methods can completely remove all knowledge related to specific entities; the Selective Unlearning setting evaluates the ability to accurately erase particular image-text associations without affecting other learned information; the Corrective relearning setting simulates a continual learning framework to examine whether previously unlearned rumors can be successfully recovered after post-training; and the Unimodal Unlearning setting assesses whether existing LLM unlearning methods can seamlessly adapt to the multimodal context of MLLMs.
Figure 1: OFFSIDE is a comprehensive benchmark for MLLMs MU, featuring four real-world settings designed to address the removal of various rumors. Texts in red represent the target rumor, while those in green indicate successful forgetting or relearning.
Five baselines are evaluated on four distinct datasets. Our comprehensive evaluation spans a variety of tasks, including classification, generation, MM-Bench (Liu et al., 2024a), and GPT evaluator. After extensive experiments, we observe several key findings, each stemming from our specially designed experimental settings, highlighting the advantages of our datasets in providing a realistic and diverse evaluation for the multimodal unlearning task. Our key contributions are as follows:
  • We propose OFFSIDE, a novel multimodal unlearning benchmark that provides four real-world scenarios (Complete Unlearning, Selective Unlearning, Corrective Relearning, and Unimodal Unlearning), demonstrating the practical value of multimodal unlearning in real-world applications.
  • OFFSIDE provides a comprehensive evaluation of unlearning, assessing forget quality, model utility, and robustness. It is the first to explore the relationship between unlearning methods and continual learning, offering valuable insights for future research.
  • After extensive experiments with five representative MU methods, we provide empirical insights into MLLM MU: (1) Unimodal unlearning methods are ineffective at unlearning multimodal rumors. (2) The unlearned rumors can be easily recovered through continual learning. (3) All baselines fail to unlearn visual rumors (rumors appear in the image). (4) All baselines are vulnerable to prompt attacks (e.g., classification task). (5) Unlearning efficacy is largely driven by catastrophic forgetting. These findings reveal key limitations of contemporary MLLM methods, thereby motivating advances specific to multimodal learning and underscoring OFFSIDE as an essential, realistic benchmark.
2 Related Work
LLM Machine Unlearning. Existing benchmarks in LLM MU have been used to test unlearning in various contexts, such as elimination of personal identification data (Patil et al., 2023), copyright protection (Eldan & Russinovich, 2023) and harmful content removal (Lu et al., 2022). Gradient Ascent (GA) (Yao et al., 2024b) was introduced to optimize the model parameters so as to maximize the removal of targeted information from the training data. However, GA often degrades performance on the retained set. Subsequent methods, including gradient descent (GD) (Liu et al., 2022), KL-based objectives (Yao et al., 2024a; Liu et al., 2024b), and “I don’t know” (IDK) losses (Maini et al., 2024), were proposed to exert finer control over the outputs of unlearned models and to mitigate collateral damage. Additionally, Negative Preference Optimization (NPO) (Zhang et al., 2024) reframes LLM unlearning as a preference-optimization problem.
MLLM Machine Unlearning.  Li et al. (2024b) introduced a token-level KL-divergence loss for MU in MLLMs, marking a pioneering attempt to apply MU in such settings. MMUNLEARNER (Huo et al., 2025) proposes a selective unlearning approach that removes visual patterns tied to a specific entity while preserving the corresponding textual knowledge within the LLM backbone. CLEAR (Dontsov et al., 2024) extends TOFU (Maini et al., 2024) by pairing personas with textual biographies and AI-generated images, and MLLMU-Bench (Liu et al., 2024c) targets the removal of private information. PULSE (Kawakami et al., 2025) extends MLLMU-Bench to include pretrained knowledge unlearning as well as continual forgetting. PEBench (Xu et al., 2025b) is the first to categorize multimodal unlearning targets into identities and events, where the targets can reside in both the textual and visual modalities. Nevertheless, the generated entities and events are overly simplistic, leading to an almost perfect unlearning effect (close to 100%), which complicates the accurate assessment of each method’s strengths and weaknesses. The benchmarks mentioned above merely evaluate the unlearned model, overlooking its potential to integrate with other post-training methods, such as continual learning. OFFSIDE is proposed to tackle the mentioned problems above: all of the images are from real-world football players, and both the image and text may contain harmful information. Additionally, we track the model’s general capabilities at different stages on MM-Bench (Liu et al., 2024a) to ensure that the unlearning process does not compromise its overall performance.
3 OFFSIDE: Unlearn football Transfer markt rumors And Relearn facts
We introduce OFFSIDE inspired by rumors in the football transfer market, where both images and accompanying text may contain inaccurate information, potentially leading the model to propagate misinformation. OFFSIDE includes 640 images of 80 football players from 20 clubs, with each image paired with 8 shared and 6 private VQA pairs. All images and their corresponding texts in our dataset are manually curated from real-world sources and cross-checked by multiple annotators, ensuring a robust benchmark for evaluating existing unlearning methods. An overview of OFFSIDE, including its data construction pipeline and evaluation procedure, is shown in Figure 2.
3.1 Models and Data Splitting
We consider a standard machine unlearning setup, with specific designs tailored for MLLMs. For all experiments, We use Qwen2.5-VL-3B and Qwen2.5-VL-7B  (Bai et al., 2025) as the base models. Let $\mathcal{D}$ denote the full dataset, which is partitioned into four disjoint subsets: $\mathcal{D}_{\textbf{forget}}$ (Forget Set), $\mathcal{D}_{\textbf{retain}}$ (Retain Set), $\mathcal{D}_{\textbf{test}}$ (Test set), and $\mathcal{D}_{\textbf{relearn}}$ (Relearn Set). In the first stage, we obtain the vanilla model by fine-tuning the pretrained MLLMs with supervision (SFT) on $\mathcal{D}_{\textbf{forget}}\cup\mathcal{D}_{\textbf{retain}}$. During the subsequent unlearning stage, various unlearning methods are applied, with access to $\mathcal{D}_{\textbf{forget}}\cup\mathcal{D}_{\textbf{retain}}$. After unlearning, we assess the model’s utility by retraining on $\mathcal{D}_{\textbf{relearn}}$ to reintroduce corrected information, simulating a continual learning setting.
The subsets can be further categorized into private and shared sets. The private set consists of QA pairs that are unique to a single image, whereas the shared set contains QA pairs that are common across multiple images of the same player. In the Selective Forget Private Information setting, we transfer the shared set of rumor-related images from $\mathcal{D}_{\textbf{forget}}$ to $\mathcal{D}_{\textbf{retain}}$, thereby simulating a more realistic scenario in which only private information is removed while shared attributes are preserved.
All four subsets are employed to enable a comprehensive evaluation. Specifically:
  • $\mathcal{D}_{\textbf{forget}}$ evaluates the effectiveness of unlearning (i.e., the extent to which the model has forgotten the targeted content);
  • $\mathcal{D}_{\textbf{retain}}$ and $\mathcal{D}_{\textbf{test}}$ assess the preservation of general model utility and knowledge (retention of non-targeted information);
  • $\mathcal{D}_{\textbf{relearn}}$ assesses the effectiveness of unlearning methods in conjunction with post-training procedures, highlighting the model’s ability to adapt in a continual learning context. In this study, we specifically examine whether the knowledge that was unlearned can be recovered through the relearning process.
The following notations distinguish different models derived from the dataset: learning algorithm $\mathcal{A}$ maps the dataset $\mathcal{D}$ to a parameterized model $\theta$ = $\mathcal{A}$($\mathcal{D}$). $\theta_{0}$ = $\mathcal{A}$($\mathcal{D}$) is the vanilla model finetuned on the full dataset. $\theta_{r}$ = $\mathcal{A}$($\mathcal{D}_{\textbf{retain}}$) denotes the retained model, which is trained from scratch on the retain set, Finally, $\theta_{u}$ refers to the unlearned model, which is produced by an unlearning algorithm $\mathcal{U}$, ideally approximating $\theta_{r}$without requiring retraining.
Figure 2: Overview of the OFFSIDE framework. The MLLM is first fine-tuned on the forget and retain set to obtain the vanilla model, during which it learns the rumors associated with each player. Various unlearning methods are then applied on forget set to obtain the unlearned model. After unlearning, the model is fine-tuned on the relearn set to simulate a continual learning setting. Performance is evaluated on four distinct subsets after both the unlearning and relearning stages.
3.2 Data Construction:
Unlike previous MLLM machine unlearning benchmarks (Dontsov et al., 2024; Liu et al., 2024c; Xu et al., 2025b), all of the data in OFFSIDE is manually curated. The data construction process consists of three stages:
Image Curation: We manually selected 80 players from 20 Premier League clubs using Google search  [^1] All images are manually selected from https://www.google.com/imghp?hl=en. For each player, we curated the following image sets: three images representing different club periods ($\mathcal{D}_{\textbf{retain}}$), one image related to a transfer rumor ($\mathcal{D}_{\textbf{forget}}$), three test images ($\mathcal{D}_{\textbf{test}}$), and one image for relearning the facts ($\mathcal{D}_{\textbf{relearn}}$). Here, $\mathcal{D}_{\textbf{test}}$ is an augmented version of $\mathcal{D}_{\textbf{retain}}$ (see Figure 2).
Text Description Curation: For each image, we constructed 14 QA pairs, comprising 6 that capture private information (e.g., the player’s market value, transfer fee) and 8 that capture shared information (e.g., the player’s height, birthdate). This design is specifically tailored for a selective unlearning setting, where the aim is to forget certain rumors or private information while retaining shared facts. Additionally, we created pure-text versions of each question-answer pair to test whether existing LLM unlearning methods can be directly extended to MLLMs.
Curating Text-Image Connections: For each player, there are 8 images representing different stages of their career at various clubs. Three images are used for training, three are reserved for testing, and one image contains a rumor about the player’s transfer history, representing the forgotten data that we aim to remove. The remaining images are used for the continual learning setting, simulating the correction of rumors. Each image is associated with 14 VQA pairs, of which 8 capture shared information such as the player’s name, nationality, and height, which are consistent across all images of the player. The remaining 6 pairs correspond to private information, specifically the player’s transfer details, with each image containing unique transfer information.
To ensure consistency across the player information and the corresponding image text, the entire dataset, covering both collection and construction, was reviewed twice by two football experts to guarantee its quality.
3.3 Evaluation Metrics
OFFSIDE provides a comprehensive evaluation framework for unlearning methods in MLLMs, assessing unlearning efficacy, generalizability, and model utility as defined by (Liu et al., 2024d), along with the model’s ability to integrate with post-training interventions (continual learning). To ensure a comprehensive evaluation, we assess the performance of the vanilla, unlearned, and relearned models on MM-Bench. We only report experimental results for each unlearning method where the model’s general capabilities are not excessively degraded. This approach guarantees that all models maintain their general capabilities throughout the process, allowing for a fair comparison of both forgetting efficacy and functional consistency.
3.3.1 Classification
To test whether the tested model can recall unlearning targets when certain rumors are provided in prompts, we design the classification task using GPT-4o to apply perturbations near the correct answer. Let $a^{n}$ represent the correct answer. Using GPT-4o, we generate a perturbation set $\mathcal{A}^{n}=\{a_{1}^{n},a_{2}^{n},a_{3}^{n},a_{4}^{n}\}$, where the model modifies the correct answer while maintaining its structure and linguistic template. Among these four responses (one correct and three perturbations), only the original is factually correct. Let ${\bm{I}}^{n}$ and ${\bm{Q}}^{n}$ denote the input images and questions, respectively, where $n$ denotes the ID of a given sample. The inputs can be denotes as (${\bm{I}}^{n}$, ${\bm{A}}^{n}$, ${\bm{Q}}^{n}$). The model predicts $\hat{y^{n}}$ by maximizing the probability $P(a^{n}\mid\displaystyle{\bm{I}}^{n},\displaystyle{\bm{A}}^{n},\displaystyle{\bm{Q}}^{n},\theta)$, where $\theta$ is the evaluated model:
$\hat{y^{n}}=\arg\max_{a^{n}_{i}\in{\displaystyle{\bm{A}}^{n}}}P(a^{n}\mid\displaystyle{\bm{I}}^{n},\displaystyle{\bm{A}}^{n},\displaystyle{\bm{Q}}^{n},\theta).$
In the unimodal setting, the input simplifies to $(\emptyset,\displaystyle{\bm{A}}^{n},\displaystyle{\bm{Q}}^{n},\theta)$. To evaluate classification performance, accuracy Acc is computed as following:
$\text{Acc}=\frac{1}{|\displaystyle{\bm{A}}^{n}|}\sum_{a^{n}_{i}\in\displaystyle{\bm{A}}^{n}}\mathbb{I}(\hat{y^{n}}=a^{n}).$
3.3.2 Generation
The generation score used in our paper is defined as the mean of the four evaluation metrics: ROUGE-1, ROUGE-2, ROUGE-L (Lin, 2004), and BLEU (Papineni et al., 2002). Specifically, it is computed as follows:
$\text{Generation Score}=\text{Mean}\left(\text{ROUGE-1}+\text{ROUGE-2}+\text{ROUGE-L}+\text{BLEU}\right).$
By averaging these four metrics, we obtain a comprehensive evaluation that captures various aspects of text generation, including lexical overlap, structural similarity, and fluency. This approach mitigates the bias of individual metrics, providing a more balanced and robust assessment of the generated content.
3.3.3 Factuality Score
Following previous work (Liu et al., 2024c), we use GPT-4o as an evaluator to assess the factuality, fluency, and semantic relevance of the generated sentences. For each question, we assign a score to the generated answer on a scale from 1 to 10. A score of 1 indicates that the content is completely incorrect or consists of meaningless symbols, while a score of 10 signifies that the answer is factually accurate and well-organized in a coherent sentence.
4 Experiment
4.1 Experiment setup
Training. We employ the Qwen2.5-VL series model as the base model for unlearning. Supervised Fine-Tuning (SFT) is performed using LoRA with a batch size of 4. For methods that require access to the retain set $\mathcal{D}_{\textbf{retain}}$, we adopt a balanced forget-retain update schedule, in contrast to the inner-loop forget and outer-loop retain strategy proposed in (Liu et al., 2024c). Specifically, we use a forget-to-retain step ratio of 1:31\!:\!3 (which corresponds to the size ratio of the forget and retain sets) to enhance training stability during the unlearning process. All experiments are conducted on a single H200 GPU (96GB).
Unlearning Algorithms. We evaluates five representative machine unlearning methods to enable an extensive analysis. Specifically, the methods examined include Gradient Ascent (GA) (Yao et al., 2024a), Gradient Difference (GD) (Liu et al., 2022), KL Minimization (Yao et al., 2024b), Preference Optimization (PO) (Maini et al., 2024), and Negative Preference Optimization (NPO) (Zhang et al., 2024). Since some of these approaches may cause progressive degradation in overall model performance during unlearning, we carefully select and report results only under conditions where the model’s core functionality is preserved, thus ensuring the practical utility of the unlearned model.
4.2 Diverse Experimental Scenarios for OFFSIDE
To better imitate complex real-world situations, we design four distinct MLLMs unlearning settings:
Complete Unlearning: In this setting, we treat each image as an individual entity, with the goal of unlearning all connections between rumor images and their corresponding text descriptions. This setting allows us to evaluate whether the unlearning algorithm can effectively forget the rumor.
Selective Unlearning: In this scenario, we focus on removing only the private information of a given image while preserving shared, non-sensitive attributes. Specifically, the shared information of $\mathcal{D}_{\textbf{forget}}$ is removed to $\mathcal{D}_{\textbf{retain}}$ and the left private information serve as the $\mathcal{D}_{\textbf{forget}}$. This approach is more realistic, as it enables the model to maintain its core ability to recognize players based on essential characteristics, such as name, height, and dominant foot.
Corrective Relearning: This setting operates within a continual learning framework, where the unlearned model, $\theta_{u}$, is allowed to relearn the facts. This not only assesses the model utility of $\theta_{u}$ but also evaluates whether the unlearned knowledge can be effectively recovered.
Unimodal Unlearning: In this setup, we combine the name of each entity with ${\bm{Q}}^{n}$. During unlearning, we set the input image to $\emptyset$. This allows us to test whether the LLM unlearning algorithms can seamlessly integrate into multimodal unlearning methods. Additionally, it aids researchers in understanding how MLLMs store knowledge.
Table 2: Results of Complete Unlearning. In this setting, we treat all 14 VQA pairs of an image as rumors. The best results of five baselines are highlighted in blue.
Models Forget Set Test Set Retain Set $M$-Bench
Class.
Acc ($\downarrow$)
Generation
Score ($\downarrow$)
Fact.
Score ($\downarrow$)
Class.
Acc ($\uparrow$)
Generation
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
Class.
Acc ($\uparrow$)
Generation
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
$M$-Bench
Acc ($\uparrow$)
Qwen2.5-VL-7B
Pretrained 49.4% 0.129 3.67 46.8% 0.115 3.66 47.2% 0.114 3.69 82.4%
Vanilla 64.4% 0.974 9.86 60.1% 0.710 5.79 65.2% 0.946 9.83 82.3%
GA 62.7% 0.616 4.97 59.0% 0.430 3.86 64.2% 0.632 5.34 81.9%
GD 23.5% 0.321 6.56 59.8% 0.521 5.05 64.3% 0.664 8.47 82.3%
KL 65.0% 0.032 0.57 60.1% 0.655 5.36 66.7% 0.861 9.20 81.9%
PO 62.9% 0.117 1.59 59.8% 0.684 5.67 64.6% 0.914 9.65 82.1%
NPO 62.1% 0.545 8.41 59.7% 0.472 5.42 64.6% 0.571 8.81 82.2%
Qwen2.5-VL-3B
Pretrained 45.5% 0.224 3.68 49.1% 0.220 3.32 49.7% 0.223 3.33 78.4%
Vanilla 53.6% 0.901 7.51 53.0% 0.651 4.67 55.3% 0.882 7.45 78.1%
GA 53.1% 0.782 6.66 52.9% 0.581 4.57 54.7% 0.774 7.30 78.0%
GD 50.5% 0.155 3.75 50.8% 0.576 4.38 53.1% 0.747 6.97 78.0%
KL 48.6% 0.550 5.62 54.1% 0.633 4.55 54.1% 0.859 7.31 78.1%
PO 57.5% 0.207 4.53 56.4% 0.671 4.00 56.4% 0.805 6.26 78.0%
NPO 45.1% 0.371 5.48 49.3% 0.337 3.71 50.2% 0.408 5.69 78.0%
Table 3: Results of Selective Unlearning. In this setting, we merely unlearn the 6 private information of the rumor images and preserve shared ones. The best results of five baselines are highlighted in blue.
Models Private Info Test Set Shared Info $M$-Bench
Class.
Acc ($\downarrow$)
Generation
Score ($\downarrow$)
Fact.
Score ($\downarrow$)
Class.
Acc ($\uparrow$)
Generation
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
Class.
Acc ($\uparrow$)
Generation
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
$M$-Bench
Acc ($\uparrow$)
Qwen2.5-VL-3B
Vanilla 56.5% 0.832 6.40 53.5% 0.654 4.66 60.8% 0.951 8.74 78.3%
GA 57.2% 0.518 5.30 52.9% 0.408 3.56 60.8% 0.709 7.52 77.9%
GD 57.3% 0.571 5.78 51.9% 0.623 4.50 60.6% 0.895 8.45 78.1%
KL 58.4% 0.725 5.20 52.2% 0.616 4.48 61.2% 0.921 8.67 78.0%
PO 59.6% 0.412 2.85 56.8% 0.545 4.02 63.7% 0.841 7.97 78.2%
NPO 58.9% 0.648 5.65 50.6% 0.584 4.29 58.9% 0.874 8.24 78.1%
Table 4: Results of Corrective relearning. In this setting, we fine-tune the unlearned model in Table 2 on $\mathcal{D}_{\textbf{relearn}}$. The best results of five baselines are highlighted in blue. The results of vanilla model directly skip the unlearning stage and relearn the facts.
Models Forget Set Test Set Retain Set Relearn Set $M$-Bench
Class.
Acc ($\downarrow$)
Generation
Score ($\downarrow$)
Fact.
Score ($\downarrow$)
Class.
Acc ($\uparrow$)
Generation
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
Class.
Acc ($\uparrow$)
Generation
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
Class.
Acc ($\uparrow$)
Generation
Score ($\uparrow$)
Fact.
Score ($\uparrow$)
$M$-Bench
Acc ($\uparrow$)
Qwen2.5-VL-7B
Vanilla 59.7% 0.576 8.36 58.3% 0.445 5.24 62.8% 0.548 8.05 59.1% 0.911 9.26 82.3%
GA 57.5% 0.584 8.49 54.4% 0.440 5.16 59.4% 0.554 8.33 55.9% 0.895 9.22 81.9%
GD 62.2% 0.489 7.87 61.4% 0.473 5.24 63.9% 0.569 8.04 62.2% 0.908 9.23 82.0%
KL 63.8% 0.336 4.55 61.5% 0.483 5.25 66.7% 0.594 8.29 62.1% 0.911 9.19 81.9%
PO 64.7% 0.567 8.23 61.2% 0.437 5.17 65.9% 0.538 7.99 65.0% 0.914 9.22 82.1%
NPO 59.3% 0.527 7.95 54.9% 0.408 5.07 59.9% 0.503 7.75 55.6% 0.909 9.21 81.9%
Qwen2.5-VL-3B
Vanilla 53.2% 0.589 6.73 53.3% 0.443 4.48 55.3% 0.522 6.39 55.2% 0.899 8.90 78.2%
GA 51.3% 0.549 6.60 52.7% 0.431 4.46 52.6% 0.501 6.21 55.1% 0.901 8.86 77.9%
GD 47.9% 0.447 6.08 47.4% 0.430 4.32 49.8% 0.505 6.11 46.7% 0.893 8.79 78.0%
KL 47.6% 0.497 6.14 48.8% 0.429 4.31 50.2% 0.512 6.38 49.2% 0.899 8.86 78.0%
PO 46.9% 0.566 6.57 47.8% 0.456 4.33 50.1% 0.501 6.41 48.8% 0.906 9.01 78.1%
NPO 47.6% 0.497 6.15 48.8% 0.429 4.31 50.2% 0.511 6.37 49.2% 0.899 8.86 77.9%
4.3 Experimental Results
In this section, we present a comprehensive comparison of several representative unlearning algorithms, evaluated using the proposed OFFSIDE across four real-world settings.
Table 2 shows the results of the setting of Complete Unlearning which is a common setting in previous works (Liu et al., 2024c; Dontsov et al., 2024). From this table, GA and NPO results in a significant drop in accuracy on both the test set and retain set while performing the forgetting process. KL and PO demonstrate strong performance on both of the Qwen2.5-VL 7B and 3B models, especially on preventing significant degradation in model performance. However, they are prone to overfitting to the $\mathcal{D}_{\textbf{relearn}}$ in practical. As a result, the methods require very careful control of the training process, limiting their practicality.
Table 3 presents the results of Selective Unlearning. We observe that all the baselines exhibit a performance drop(compared to the vanilla model) in both private information and shared information. This indicates that the tested baselines have trouble selectively unlearning private information in a given image while preserving shared information. This uncovers that existing methods focus on entity-level unlearning, which disrupts all associations between a given image and related text, making it challenging to selectively cut part of the associations.
Table 4 presents the results of Corrective Relearning. The model used here is based on Table 2, where we retrain the unlearned model. Surprisingly, we found that after relearning, all of the baselines exhibit a "bounce-back" effect on either the 3B or 7B model (since relearning serves as a continual learning process, it is common to forget, but uncommon to recover), indicating that the knowledge previously forgotten can be easily recovered through simple retraining. Specifically, KL achieves a fact score of 0.57 on the forget set, which increases to 4.55 after relearning. This suggests that none of the baselines truly forget the rumor information; instead, they merely conceal it.
Figure 3 presents the results of the Unimodal Unlearning setting. In the multimodal setup, the input consists of both text and images, while in the unimodal setup, only text is provided. As shown in the results, multimodal unlearning methods, while forgetting information, simultaneously lead to a trade-off with a decrease in model utility. All unimodal unlearning methods struggle to unlearn multimodal rumors. This suggests that the target information is not only restored in LLMs but also embedded within the visual layer of MLLMs. This highlights the need for researchers to design unlearning methods specifically tailored to the unique characteristics of MLLMs.
Figure 3: Results of the Unimodal Unlearning. RS, TS, FS represent retain set, test set, and forget set, respectively. CA, GS, FS refer to classification accuracy, generation score, and fact score, respectively.
4.4 Discussion
In this section, we present and discuss several key findings based on the experimental results, and we summarize the main conclusions drawn from our analysis.
Figure 4: Illustration of experimental conclusions, observed from the OFFSIDE benchmark.
All baselines have trouble unlearning visual rumors. As shown in Figure 4, when confronted with deceptive visual rumors, the model can easily be misled due to its powerful reasoning capabilities. This is intuitive because, even if the model forgets the visual rumors at the visual-text fusion level, it still lacks the knowledge needed to answer the question. Consequently, the model’s response largely depends on the information it perceives in the image, without recognizing that the visual information is unreliable.
All of the tested baselines remain vulnerable to prompt based attacks. Although certain methods achieve low generation and fact scores on the forget set, they still maintain high classification accuracy. This indicates that when rumor information appears in the prompt, the model can still recognize and select the incorrect knowledge, thereby exposing its susceptibility to prompt-induced retrieval. For instance, as shown in Table 2, PO demonstrates strong performance in generation and fact scoring, suggesting effective forgetting. However, its classification accuracy remains close to that of the original, unmodified model, revealing a critical gap in current unlearning approaches. This persistent ability to match forgotten content in classification task underscores the need for more comprehensive and robust unlearning techniques that address both generative and discriminative exposure.
Unlearning efficacy is largely driven by catastrophic forgetting. In Figure 4, we compare the GPT-evaluation results of models relearned after forgetting with those of the directly relearned vanilla model. We observe that the knowledge unlearned by the baselines closely resembles catastrophic forgetting in continual learning. Specifically, the unlearned sample IDs through GA, GD, KL, and NPO show 71%, 48%, 58%, and 60% similarity to the forgotten IDs after a simple relearning step. This suggests that the unlearning ability of the tested baselines is primarily driven by catastrophic forgetting in continual learning. This is largely because these unlearning algorithms can be viewed as a form of continual learning, which inherently results in catastrophic forgetting. This phenomenon demonstrates how catastrophic forgetting can be leveraged as a method for machine unlearning and highlights a promising direction for future research.
In some rare cases, the unlearned model outperforms the vanilla model. As illustrated by the PO example in Table 2, the unlearned model achieves a higher generation score on the test set compared to the vanilla model. This improvement can be primarily attributed to the reintroduction of $\mathcal{D}_{\text{retain}}$. To obtain the vanilla model, we ensure that it is not overfitted to $\mathcal{D}_{\text{finetune}}$. During the unlearning process, incorporating $\mathcal{D}_{\text{retain}}$ can enhance generalization on $\mathcal{D}_{\text{finetune}}$. However, methods that rely on $\mathcal{D}_{\text{retain}}$ are at risk of overfitting, which requires careful management.
Methods such as KL Minimization demonstrate greater effectiveness when applied to a 7B model, but show reduced efficacy with a 3B model. This is primarily due to the random direction of optimization in gradient-ascent-based methods. Before model collapse occurs, these methods struggle to control the optimization direction, which may lead to significant deviations in the results. In contrast, methods like PO, which do not rely on gradient ascent, show more stable performance across both models.
5 Conclusion
We introduce OFFSIDE, designed to simulate a real-world scenario for unlearning in MLLMs. We propose four distinct settings (Complete Unlearning, Selective Unlearning, Corrective Relearning, and Unimodal Unlearning) to establish a robust unlearning framework and comprehensively evaluate a list of representative machine unlearning baselines. Our findings indicate that: all baselines struggle to unlearn visual rumors, and the unlearned knowledge can be easily recovered through prompt attacks (classification tasks) or simple relearning. Moreover, directly applying unimodal unlearning methods proves inadequate for effectively removing multimodal rumors, highlighting the need for algorithms specifically tailored to MLLM unlearning. Notably, our corrective relearn setting reveals that the unlearning ability of the tested baselines is primarily driven by catastrophic forgetting within continual learning. This suggests a potential connection between machine unlearning and continual learning. Overall, our findings provide valuable empirical insights that guide the development of more effective unlearning methods for future MLLM research.
Ethics statement.
All images used in this study are collected from real-world sources. To prevent the spread of misinformation or rumors, we will provide the original URLs for each image. These URLs are intended solely for research purposes and should be used exclusively by qualified researchers for academic investigation. This transparency ensures traceability and supports reproducibility while upholding responsible research practices.
Reproducibility statement.
The computational resources and experimental setup required for this study are provided. Upon acceptance of the paper, we will release all source code associated with our experiments.
Appendix
The Appendix is organized as follows.
  • Section A: Introduces the baseline methods compared in our experiments.
  • Section B: Introduces the MM-Bench Indicator Definitions.
  • Section C: Describes the supervised Fine-tuning process.
  • Section D: A detailed version of hyperparameters settings.
  • Section E: A case study of our proposed settings.
  • Section F: Outlines the extent to which we employ LLMs.
  • Section G: A detailed description of GPT prompt strategy.
Appendix A Unlearning Methods
Gradient Ascent(GA) Yao et al. (2024b): This method updates the model parameters by maximizing the likelihood of misprediction for the samples in the forget set $D_{\text{forget}}$. For a given sample $x\in D_{\text{forget}}$, the loss function is defined as:
$\mathcal{L}(D_{\text{forget}},w)=\frac{1}{|D_{\text{forget}}|}\sum_{x\in D_{\text{forget}}}\ell(x,w).$ (1)
Gradient Difference (GD) Liu et al. (2022): This method extends gradient ascent by simultaneously focusing on forgetting the samples in the forget set $D_{\text{forget}}$ while preserving performance on the retain set $D_{\text{retain}}$. The objective is to balance increasing the loss on the forget set and minimizing its impact on the retain set. The overall loss function to be minimized is formulated as:
$\mathcal{L}_{\text{diff}}(w)=-\mathcal{L}(D_{\text{forget}},w)+\mathcal{L}(D_{\text{retain}},w).$ (2)
KL_Min Yao et al. (2024a): This method extends gradient ascent by introducing an additional objective that minimizes the Kullback–Leibler (KL) divergence between the predictions of the original model $M_{\text{ori}}$ and the updated model $M_{\text{new}}$ on the retain set $D_{\text{retain}}$. The KL divergence loss is defined as:
$\mathcal{L}_{\text{KL}}=\frac{1}{|D_{\text{retain}}|}\sum_{s\in D_{\text{retain}}}\frac{1}{|s|}\sum_{i=2}^{|s|}\mathrm{KL}\!\left(M_{\text{ori}}(s_{<i})\;\Big\|\;M_{\text{new}}(s_{<i})\right).$ (3)
The overall training objective combines the gradient ascent loss on the forget set with the KL divergence loss on the retain set, which is formulated as:
$\mathcal{L}_{\text{total}}(w)=-\mathcal{L}(D_{\text{forget}},w)+\mathcal{L}_{\text{KL}}.$ (4)
Preference Optimization (PO) Maini et al. (2024): This method steers the model to align with newly generated responses such as “I do not know the answer” and its variants for questions belonging to the forget set $D_{\text{forget}}$. At the same time, it incorporates a retain-set term to ensure that predictions on the retain set $D_{\text{retain}}$ remain unaffected. The total objective function is formulated as:
$\mathcal{L}_{\text{idk}}(w)=\mathcal{L}(D_{\text{retain}},w)+\mathcal{L}(D_{\text{forget}}^{\text{idk}},w).$ (5)
Negative Preference OptimizationZhang et al. (2024): In our work, we adopt the Negative Preference Optimization (NPO) technique to unlearn undesirable data, thereby mitigating the catastrophic collapse often observed in gradient ascent–based methods. NPO builds on the preference optimization framework, but specifically targets negative samples from the forget set $D_{\text{forget}}$.
The NPO loss is defined as:
$\mathcal{L}_{\text{NPO}}=\frac{2}{\beta}\;\mathbb{E}_{(x,y)\in D_{\text{forget}}}\Bigg[\log\!\left(1+\Big(\tfrac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\Big)^{\beta}\right)\Bigg],$ (6)
where $\pi_{\theta}(y|x)$ denotes the probability assigned by the current model, and $\pi_{\text{ref}}(y|x)$ is the probability from a reference model trained on the entire dataset. The parameter $\beta$ controls the smoothness of optimization: as $\beta\to 0$, the NPO loss converges to the standard gradient ascent loss.
By minimizing $\mathcal{L}_{\text{NPO}}$, the model reduces its reliance on the forget set, leading to a more stable unlearning process and avoiding the rapid degradation characteristic of gradient ascent. In our experiments, we follow the original paper and set $\beta=0.9$. The reference distribution $\pi_{\text{ref}}$ is obtained by fine-tuning the pre-trained model exclusively on the retain set $D_{\text{retain}}$.
Table 5: Performance of the vanilla OFFSIDE and MLLMU-Bench models on MM-Bench.
Method $M$-Bench
Overall LR AR $R$ FP-S FP-C CP
Qwen2.5-VL-7B 82.4 71.7 84.9 80.2 89.8 80.1 81.3
LLaVA-1.5-7B 62.3 29.9 73.1 54.7 69.6 57.7 68.5
MLLMMU-Qwen2.5-VL-7B 80.4 68.2 80.2 73.9 87.9 77.7 83.2
OFFSIDE-Qwen2.5-VL-7B 82.3 69.2 82.0 79.1 88.5 78.9 85.5
Appendix B MM-Bench Indicator Definitions
To comprehensively evaluate model capabilities, MM-Bench defines multiple indicators that jointly cover overall performance, reasoning ability (attributes and relations), and perception ability at both fine-grained and coarse-grained levels. These indicators aim to capture the model’s strengths and weaknesses across diverse dimensions of multimodal understanding.
Overall: Overall denotes the overall accuracy of a model on the entire MM-Bench-test set. It reflects the model’s performance across all ability dimensions, encompassing both perception and reasoning tasks, and is evaluated under the strict circularEval strategy.
Attribute Reasoning(AR): AR measures a model’s ability to reason about attributes of objects or people. This includes identifying physical properties such as hardness or conductivity, inferring the function of tools and objects, and recognizing identities or professions based on appearance.
Relation Reasoning(RR): RR measures reasoning about different types of relationships. It includes social relations between people (e.g., family, friends, colleagues), physical relations in the environment (such as spatial positioning or distance), and natural relations in ecosystems (such as predation, competition, or symbiosis).
Fine-grained Perception(FP-S): FP-S reflects the model’s fine-grained perception ability when dealing with a single object or entity. It covers tasks such as locating objects in an image, recognizing specific attributes like shape or color, identifying celebrities or famous figures, and reading text within an image (OCR).
Fine-grained Perception(FP-C): FP-C measures fine-grained perception across multiple objects in an image. It includes understanding spatial relationships between objects, comparing attributes (e.g., colors or shapes), and recognizing human actions and interactions involving multiple participants.
Coarse Perception(CP): CP evaluates coarse-grained perception abilities. It focuses on a model’s capacity to recognize general aspects of an image, such as its style (photo, sketch, painting), the scene it depicts (indoor, forest, street), the overall emotion it conveys (happy, sad, anxious), the visual quality (clarity, brightness, contrast), and the main topic or subject.
In Table 5, we use MLLMMU-Bench and OFFSIDE to fine-tune Qwen2.5-VL 7B with the same number of steps. We find that fine-tuning on synthetic datasets reduces the model’s general ability. However, using the proposed OFFSIDE method preserves the model’s general performance. This highlights the importance of using a dataset that simulates real-world scenarios.
Appendix C Vanilla Model Fine-tuning
To simulate a real-life scenario where unlearning algorithms are applied to a “pre-trained" model, we first fine-tune the off-the-shelf MLLM model using the full set $\mathcal{D}$. The fine-tuning process involves pairing visual inputs (images of the individuals) with textual information (questions and answers), allowing the model to learn associations between these modalities. For each input $\langle{\bm{I}}^{n},{\bm{Q}}^{n},{\bm{Y}}^{n}\rangle$, where ${\bm{I}}^{n}$ is the visual representation of the individual, ${\bm{Q}}^{n}$ is the question, and ${\bm{Y}}^{n}$ is the ground-truth answer, the model is trained to predict the answer $\hat{y^{n}}$. The loss function for a single sample is defined as the negative log-likelihood (NLL) over the answer tokens:
$l({\bm{Q}}^{n},{\bm{Y}}^{n},w)=\frac{1}{|{\bm{Y}}^{n}|}\sum_{i=1}^{|{\bm{Y}}^{n}|}\text{NLL}_{w}(y_{i}^{n}|[{\bm{Q}}^{n},y_{<i}^{n},{\bm{I}}^{n}]),$
where $w$ represents the model parameters, and the loss is averaged over all tokens in the answer sequence ${\bm{Y}}^{n}$. The overall objective during fine-tuning is to minimize the average loss across the entire dataset $D$, expressed as:
$L(D,w)=\frac{1}{|D|}\sum_{({\bm{Q}}^{n},{\bm{Y}}^{n})\in D}l({\bm{Q}}^{n},{\bm{Y}}^{n},w).$
After fine-tuning, the model represents the "vanilla" version, which serves as the starting point for subsequent unlearning experiments.
Appendix D Hyperparameters Settings
For all fine-tuning phases, we set the maximum output length to 128. For the LoRA configuration, we set $r=8$, $\alpha=32$, dropout = 0.05, and the learning rate to $1\times 10^{-4}$. For unlearning methods, we maintain the same settings except for the learning rate, which is adjusted to $2\times 10^{-5}$. For methods requiring $\mathcal{D}_{\textbf{retain}}$, the previous benchmark utilized an inner loop for the forget set and an outer loop for the retain set. This setup meant that the impact of the forget loss could be easily "healed" by gradient descent on retain batches, which introduced significant randomness due to the instability of the tuning process. To address this issue, we adopted a balanced forget-retain update strategy (e.g., forget step: retrain step = 1:3), ensuring more stable and consistent results. We will provide more detailed Hyperparameters setting in our code.
Figure 5: Case study of four unlearning settings, each simulating a real-world MLLM unlearning scenario.
Appendix E Case study
We present the case study under our specially designed four settings in Figure 5. Complete Unlearning evaluates the ability of MU methods to remove all image-text connections, ensuring that the model forgets the entire knowledge associated with specific visual or textual inputs. Selective Unlearning tests the methods’ capacity to accurately unlearn unwanted knowledge while preserving the shared, valuable information across modalities, highlighting the precision of the unlearning process. Relearn Facts serves as a continual learning setting, where the model must relearn certain facts after unlearning them, simulating real-world scenarios where knowledge evolves and needs to be updated. Finally, Unimodal Unlearning examines whether unimodal methods, designed for single-modality data, can be directly applied to Multimodal Large Language Model (MLLM) MU settings, revealing the limitations and challenges of using unimodal techniques in multimodal contexts.
Appendix F Use of AI Assistants
LLMs are employed to polish the language of our paper. What’s more, we evaluate the factual accuracy of the generated answers using GPT-4o. Apart from these, we have not included any usage of LLMs, preserving the originality and quality of this work.
Appendix G GPT Prompt Strategy
In this section, we detail the methodology employed to construct our dataset using the OpenAI API. To evaluate the faculty score of the generated answers, we carefully designed a structured prompt, as illustrated in Figure 8. This prompt enables a systematic and transparent evaluation of generated answers by providing clear, multi-dimensional criteria focused on factuality, relevance, and fluency. It ensures consistency and granularity through a well-defined scoring scale and explicit guidelines for handling language issues. Furthermore, we leverage GPT-4o to generate high-quality classification data, with the exact prompt used provided in Figure 7. In addition to classification data, we also utilize GPT-4o to construct unimodal unlearning data, as detailed in the prompt shown in Figure 6. This type of data is specifically designed to isolate and examine individual modalities or attributes within the model’s knowledge.
Figure 6: Prompt strategy of creating pure text description.
Figure 7: Prompt strategy of creating classification data.
Figure 8: Prompt strategy of evaluating factuality score through GPT-4o.